OpenStudy (anonymous):

Which of the following is the rational exponent expression of third root of 4 n?

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

(4n)3 3n4 quantity of 4 n to the one third power 4 times n to the one third power

OpenStudy (anonymous):

@DebbieG

OpenStudy (anonymous):

@Psymon

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

(4n)3 ? @jim_thompson5910

OpenStudy (anonymous):

4^ n/3 @jim_thompson5910 ?

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (akashdeepdeb):

Square root of anything = Itself powered by 1/2 Cube root of anything = power by 1/3 Fourth root of anything = powered by 1/4 Now what do you think the third root of 4n would be? :)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (akashdeepdeb):

@Nurialozza96 ? :/

OpenStudy (anonymous):

@AkashdeepDeb im lostt /:

OpenStudy (akashdeepdeb):

Square root of 2 can also be written as (2)^1/2 Cube root or third root of any number n is (n)^1/3 So what will the third root of any number 4n be? :)

OpenStudy (zzr0ck3r):

all you need is \[a^{\frac{b}{c}}=\sqrt[c]{a^b}\] you have\[\sqrt[3]{4^n}\]or\[\sqrt[3]{4n}=\sqrt[3]{(4n)^1}\]I cant tell because you just put 4 n

OpenStudy (anonymous):

Which of the following is the rational exponent expression of third root of 4 n? thats the question .. choice are ... (4n)3 3n4 quantity of 4 n to the one third power 4 times n to the one third power

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

i think its A? @zzr0ck3r

OpenStudy (zzr0ck3r):

we have told you more than enough to answer this question, if you dont understand then ask a question. Dont just say you dont understand. ask a specific question. If i tell you the answer you will not be able to do it when you are taking a test....

OpenStudy (anonymous):

I did ask questions. Im just lost about the WHOLE THING in genera. math isnt my thing and as much as I try to process it, it just doesnt go through

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

does your question mean 4n or 4^n?

OpenStudy (anonymous):

4n

OpenStudy (zzr0ck3r):

ok so you want the 3rd root of 4n notice that (4n) = (4n)^1 so you want the 3rd root of 4n to the first power does this make sense so far?

OpenStudy (zzr0ck3r):

and please stop tagging me every 3 mins:)

OpenStudy (zzr0ck3r):

hello?

OpenStudy (anonymous):

yes that makes sense my apologies my comp died

OpenStudy (zzr0ck3r):

\[(4n)^{\frac{1}{3}}\]